OpenStudy (anonymous):

find the absolute minimum value of the function f(x) = 1/4x^4 + x^3 - 5x^2

OpenStudy (shamil98):

Are you allowed to use calculus? or is this for algebra 2?

OpenStudy (anonymous):

This is calculus

OpenStudy (shamil98):

You start by taking the first derivative right?

OpenStudy (anonymous):

Yes, I took the first derivative but after that I'm not sure. I can set it to zero but I'll have an x in all my numbers. I could factor out the x but after doing that and then setting it to zero, I don't get the right answer

OpenStudy (shamil98):

f(x) = 1/4x^4 + x^3 - 5x^2 f'(x) = x^3 + 3x^2 - 10x 0 = x^3 + 3x^2 - 10x 0 = x(x^2 + 3x - 10) 0 = (x+5)(x-2) can you do it from here?..

OpenStudy (shamil98):

btw 1/4x^4 derived is 4/4x^3 which is just x^3, if you're wondering

OpenStudy (anonymous):

That is what I did initially however the answer in my book shows -93.75

OpenStudy (shamil98):

your book isn't wrong.

OpenStudy (shamil98):

(x+5)(x-2) = 0 x = -5 , x = 2 evaluate x at -5 f(x) = 1/4x^4 + x^3 - 5x^2 you get -93.75

OpenStudy (anonymous):

ahhhhhhhhhh, yes i see it now. Thank you very much!

OpenStudy (shamil98):

Glad to be of help.